OpenStudy (vishweshshrimali5):

Evaluate: \[\large{1^2. 2! + 2^2. 3! +... + n^2 n!}\]

OpenStudy (vishweshshrimali5):

@Mertsj @hartnn @UnkleRhaukus @Callisto

OpenStudy (vishweshshrimali5):

\[\sum_{1}^{n}(n^2.(n+1)!)\]

OpenStudy (vishweshshrimali5):

This looks better.

OpenStudy (anonymous):

\[\sum_{k=1}^nk^2(k+1)!\]

OpenStudy (vishweshshrimali5):

\[\sum_{1}^{n}(k^2.(k+1)!)\] \[= \sum_{1}^{n}(k^2.(k+1)k (k-1)!)\] \[=\sum_{1}^{n}(k^2.(k+1).(k-1)!\]

OpenStudy (vishweshshrimali5):

*k^3 (in last step

OpenStudy (vishweshshrimali5):

\[\sum_{1}^{n}(k^3.(k+1).(k-1)!)\]

OpenStudy (vishweshshrimali5):

\[\sum_{1}^{n}(k^4 .(k-1)!) + \sum_{1}^{n}(k^3 . (k-1)!)\]

OpenStudy (vishweshshrimali5):

No idea after that........... :(

OpenStudy (anonymous):

I think you'll have to use the third one from the bottom in the list: http://en.wikipedia.org/wiki/Summation#Some_summations_involving_binomial_coefficients_and_factorials \[\sum_{k=0}^nk\cdot k!=(n+1)!-1\] Wolfram gives a result that looks a lot like it: http://www.wolframalpha.com/input/?i=sum+k%5E2%28k%2B1%29%21+from+k%3D1+to+n

OpenStudy (vishweshshrimali5):

But, here we have (k+1)!k^2

OpenStudy (anonymous):

You'll have to do some rearranging. I'm not sure how to go about this myself...

OpenStudy (vishweshshrimali5):

Okk I will try

OpenStudy (vishweshshrimali5):

Thanks

mathslover (mathslover):

No on?